OpenStudy (anonymous):

Proove that two Fermat numbers always are relatively prime. I have the lead that F_m | (F_n - 2) for m

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

I've tried to understand the solution here: https://www.artofproblemsolving.com/Wiki/index.php/Fermat_number but don't really understand it.

OpenStudy (anonymous):

I also have this solution but don't understand it neither

myininaya (myininaya):

Hey, so from that link the only thing I don't understand so far is why all of that shows that they are relatively prime. I do understand everything else though.... So still thinking as to why it shows they are....

OpenStudy (anonymous):

I thinks it has to do with the fact that every Fermat number is odd so can not be divided by two and hence the only gcd is 1

myininaya (myininaya):

Yeah. lol. I think you are right.

OpenStudy (anonymous):

Woho! :)

OpenStudy (anonymous):

But who do they get the recursion formula from \[F _{n}=2^{2^{n}}+1\]

myininaya (myininaya):

The induction part confuses you?

myininaya (myininaya):

\[F_0=2^{2^0}+1=2^1+1=2+1=3 \] \[F_1=2^{2^1}+1=2^2+1=4+1=5=3+2 \] \[F_2=2^{2^2}+1=2^4+1=16+1=17=15+2 =3(5)+2=F_0 \cdot F_1 +2 \]

OpenStudy (anonymous):

Wonderful!

OpenStudy (anonymous):

So when one have the recursion formula \[F _{n+1}=F _{0}*F _{1}*F _{2}*...*F _{n}+2\] The solution is to show that the recursion formula is correct using induction, right?

myininaya (myininaya):

Yes :)

myininaya (myininaya):

They are proving it :) Proving that will prove gcd(Fm,Fn)=1

myininaya (myininaya):

If you need help with the algebra, I can do that. Right now, I'm multitasking though...so I will help slowly.

OpenStudy (anonymous):

So there are two things I don't understand the first is:\[F _{k+1}=2^{2^{k+1}}+1\] \[=\left( \left( 2^{2^{k}} \right)^{2} +2*2^{2^{k}}+1 \right) - 2*2^{2^{k}}\]

OpenStudy (anonymous):

The second thing is \[\gcd(F _{m},F _{n})=\gcd(F _{m},2)\]

myininaya (myininaya):

\[=2^{2^{k+1}}+1=2^{2^k2^1}+1=2^{2^k 2}+1=(2^{2^k})^2+1\]

myininaya (myininaya):

They added in a zero.

OpenStudy (anonymous):

Why did the add the zero for?

myininaya (myininaya):

\[=(2^{2^k}+1)^2+1-2 \cdot 2^{2^k} \] So they can write that one part as Fk

myininaya (myininaya):

or F_0F_1F_2...F_(K-1)+2

myininaya (myininaya):

They were setting up for the induction part

myininaya (myininaya):

\[=(2^{2^k}+1)(2^{2^k}+1)-2^{2^k}\] Ignore that 1oops

myininaya (myininaya):

the one in expression i wrote before this one

myininaya (myininaya):

We are trying to show F_0F_1....F_k+2

OpenStudy (anonymous):

Ok I think I get the whole induction proof now.

OpenStudy (anonymous):

So seeing that \[F _{k+1}=F _{0}*F _{1}...*F _{k-1}*F _{k}+2\] How do I know that \[GCD(F _{m},F _{n})=GCD(F _{m},2)\]

myininaya (myininaya):

Please ask any questions if you have them. I like this proof.

OpenStudy (anonymous):

* given that m<n

myininaya (myininaya):

Well gcd(odd,2)=1 since odd aren't divisible by 2

OpenStudy (anonymous):

Sorry formulated it bad, I mean how do I know that gcd F_n is 2 ?

myininaya (myininaya):

no Fn is not equal to 2

OpenStudy (anonymous):

* F_n is replaces by two *

OpenStudy (anonymous):

So where does the replacement of F_n to 2 come from?

myininaya (myininaya):

gcd(Fm,Fn) = gcd(Fm,F0F1...F(n-1)+2) What about this?

OpenStudy (anonymous):

Sorry don't understand, Isn't that equal to f_(n+1)?

OpenStudy (anonymous):

Has it to do with the initial lead F_m | (F_n - 2) ?

myininaya (myininaya):

Well F_(n+1) equals F0F1....Fn+2 ---- Anyways, Fm|(Fn-2) =>for some integer k we have k*Fm=Fn-2 => k*Fm+2=Fn So gcd(Fm,Fn) = gcd(Fm, k*Fm+2) Let gcd(Fm, k*Fm+2)=d => for some integers a and b we have aFm+b(k*Fm+2)=d Fm(a+bk)+2b=d But a+bk and b are just integers We know gcd(Fm,2)=1 since Fm is odd and isn't divisible by 2. This implies for some integers s and t we have sFm+2t=1 where s above is a+bk and t is b from above. Did this make it more confusing?

myininaya (myininaya):

Implies d=1

OpenStudy (anonymous):

Not at all, that made perfect sense actually

OpenStudy (anonymous):

Will have to put it all together and solve it 2-3 times just to make sure it stays in my brain but you're given me the broad understanding, can't thank you enough!

myininaya (myininaya):

I try.

OpenStudy (anonymous):

Thank you for your patience, have a great evening!

myininaya (myininaya):

You too. :)